OpenStudy (anonymous):

SHOW ALL OF THE STEP THAT YOU USE TO SOLVE THE PROBLEM 19. GI bisects

OpenStudy (chris911):

then those two angles are congruent, x-3 = 2x-13 just solve for x from here....

OpenStudy (anonymous):

x−3=2x−13 x−2x=−13+3 −x=−10 x=10